2 salesmen, A and B, working in a certain district participated in a survey


conducted by the Head office. It was found that, number of sales made by


A and B are 20 and 18 respectively. Average sales made by A was of about


Rs 170 while for B it was Rs. 205. Standard deviation deviation for A is 20


and for B it is 25. Test whether there is any significant difference in the


average sales between the 2 salesmen at 5% level of significance.

1
Expert's answer
2022-02-24T02:28:02-0500

Difference in sales = 20-18 = 2, difference in deviation = 25-20 =5. The total difference is 7. Average of the lower values = (18+20)/2= 19

7/19 = 0.36. This clearly indicates there is significant difference in the average sales.